### 01 Feb 2024 v1

# Real-Time Coordination of Integrated Transmission and Distribution Systems: Flexibility Modeling and Distributed NMPC Scheduling

Abstract
This paper proposes a real-time distributed operational architecture to efficiently coordinate integrated transmission and distribution systems (ITD). At the distribution system level, the distribution system operator (DSO)computes the aggregated flexibility of all controllable devices by power-energy envelopes and provides them to the transmission system operator (TSO). At the transmission system level, a distributed nonlinear model predictive control (nmpc)approach is proposed to coordinate the economic dispatch of multiple TSOs, considering the aggregated flexibility of all distribution systems. The subproblems of the proposed approach are associated with different TSOs and individual time periods. In addition, the aggregated flexibility of controllable devices in distribution networks is encapsulated, re-calculated, and communicated through the power-energy envelopes, significantly reducing computational costs and eliminating redundant information exchanges between TSOs and DSOs for privacy and security preservation. The extensive numerical experiment demonstrates the scalability and the computational efficiency of the proposed distributed operational architecture.
Index Terms:
Data Preservation, Distributed Nonlinear Model Predictive Control, Flexibility Aggregation, Integrated Transmission and Distribution Systems, Multiperiod AC Optimal Power Flow.
[…]
This work was supported in part by the BMBF-project ENSURE II with grant number 03SFK1F0-2, in part by the Swiss National Science Foundation (SNSF) under the NCCR Automation project, grant agreement 51NF40_180545, and in part by the Swiss Federal Office of Energy’s “SWEET” programme and performed in the PATHFNDR consortium.
I Introduction
With the rapid adoption of distributed energy resources (DERs)in distribution systems, the aggregated flexibility of all these controllable devices can play an important role in dispatch problems in transmission systems. It can improve the operational efficiency of the overall power grid and enhance reliability when integrating increased levels of renewable energy resources [1]. Hence, coordinating integrated transmission and distribution systems (ITD)becomes essential for efficiently operating future power systems [2, 3].
Multiperiod dispatch problems for ITD systems usually couple individual steady-state optimal power flow (OPF)optimization problems over multiple time periods [4, 5, 6]. The coupling constraints include the generator ramping limits, the model of distributed energy storage systems (ESS), and other time-dependent constraints to consider the controllable devices with time-variant properties. However, it is still a challenge to solve a multiperiod AC optimal power flow (MPOPF)for ITD systems. On the one hand, the AC OPF is generally NP-hard [7], and the complexity of solving an MPOPF is further magnified by the intercoupling of subsequent time periods [5]. On the other hand, collecting necessary and realistic data from multiple stakeholders (i.e., TSOs and DSOs), including grid topology, load profiles, and other sensitive information regarding consumer behaviors, is either not preferred or restricted by regulations [3]. To address these challenges and achieve an efficient operation of the overall ITD, recent research analyzed the determination of the aggregated flexibility of the controllable devices in distribution networks [8]. The flexible dispatch region of a distribution network is summarized in a time-coupled power-energy band, taking into account the network topology  [9] and operational constraints. However, the proposed ITD framework does not consider the coordination between multiple TSOs in a data-preserving manner, and the proposed inner approximation is computationally inefficient, requiring solving multiple mixed-integer linear programming (milp)problems iteratively.
To enable privacy preservation and improve computational efficiency, distributed operation frameworks enable TSOs and DSOs to operate independently and collaborate effectively by sharing limited information with a subset of other operators [10, 11, 12, 13]. These proposed distributed frameworks can maintain data privacy and decision-making independence and are based on distributed AC OPF [11, 12] and MPOPF with receding horizon [13]. In addition to the aforementioned distributed algorithms, aladin is proposed for generic nonconvex optimization problems with convergence guarantees in [14]. aladin-type algorithms have been successfully applied to solve the single period AC OPF for heterogeneous power systems [15]. However, these aforementioned studies either lack a convergence guarantee or their scalability is limited by the computational complexity, which so far hinders an application to MPOPF in ITD systems.
In this paper, we propose an economic dispatch problem for ITD systems over multiple periods and utilize an aladin-type distributed nmpc to solve the optimization problem efficiently while preserving data privacy. The major contributions of this paper are summarized as follows:
  1. 1.
    We propose a novel real-time framework that combines the flexibility aggregation method [1] and distributed optimization [3] for coordinating the economic dispatch problem of ITD systems. At the distribution system level, the DSO computes the feasible dispatch region of all controllable devices leveraging the LinDistFlow model [16]. This region is communicated to the TSO. At the transmission system level, considering the aggregated flexibility of distribution systems, the TSO s solve the coordinated economic dispatch problem using a distributed approach. The scheme of the proposed operational architecture is shown in Fig. 1, as inspired by the actual situation in Germany.
  2. 2.
    In contrast to our previous work [3], we develop an aladin-type distributed nmpc approach for the multiperiod coordination of ITD systems within the proposed real-time framework. This approach is capable of decoupling the large-scale dispatch problem associated with different system operators and individual time periods. This particular design distributes the computational complexity of MPOPF over different stakeholders for computational affordability while maintaining the privacy of relevant information. Furthermore, the measurement data (i.e., load profiles and solar and wind outputs) from the ENTSO-E Transparency Platform [17] is employed in a simulation, showcasing the proposed methodology’s effectiveness with over 100,000 state variables divided into 400 subproblems, demonstrating its scalability and practical applicability in TSO-DSO coordination.
(a) Aggregated flexibility of DSOs
(b) TSOs Coordination
Figure 1: Proposed real-time coordination of integrated transmission and distribution systems
II Problem Formulation
This section presents a distributed framework for coordinating ITD systems. As shown in Fig. 1, at the distribution level, each DSO calculates its own feasible region taking into account its controllable devices and provides it to the corresponding TSO. The TSO then solves a coordinated economic dispatch problem over multiple time periods, considering the aggregated flexibility of the DSOs. Throughout this paper, solar and wind generation are considered as negative demands.
II-A Model of Flexibility in Distribution Systems
In this section, we consider a radial distribution system denoted by a directed tree graph $\mathcal{G}(\mathcal{N},\mathcal{L})$, where $\mathcal{N}=\{1,...,N_{\mathrm{bus}}\}$ is the set of buses. The set $\mathcal{L}\subseteq\mathcal{N}\times\mathcal{N}$ collects “links” or “lines” for all $(i,j)\in\mathcal{L}$. The number of links in a distribution network is $N_{\textrm{line}}$. Bus $1$ is the slack (root) bus and is assumed to have a fixed voltage. We also assume that the distribution systems have a pure tree topology, i.e., $N_{\textrm{bus}}=N_{\textrm{line}}+1$ holds. We leverage the definition of connectivity matrices $C^{g}$, $C^{s}$ and $C^{\textrm{pcc}}$ with respect to generator, ESS and the point of common coupling (PCC)between transmission and distribution, as defined in [18].
Definition 1 ([19]).
[…]
II-A1 Exact Feasible Set
We use the LinDistFlow model [9] to describe the relationship between the voltages and net loads in distribution systems by the following linear power flow equation:
$1\,=\,$ $e_{1}^{\top}U_{k},$ (1a)
$0\,=\,$ $C^{\textrm{inc}}U_{k}-2RP^{l}_{k}-2XQ^{l}_{k},$ (1b)
$0\,=\,$ $e_{1}p^{\textrm{pcc}}_{k}-P^{d}_{k}-(C^{\textrm{inc}})^{\top}P^{l}_{k}-C^{s}P^{s}_{k},$ (1c)
$0\,=\,$ $e_{1}q^{\textrm{pcc}}_{k}-Q^{d}_{k}-(C^{\textrm{inc}})^{\top}Q^{l}_{k},$ (1d)
$\underline{U}\,\leq\,$ $U_{k}\leq\overline{U},$ (1e)
$\underline{P}^{s}\,\leq\,$ $P^{s}_{k}\leq\overline{P}^{s},$ (1f)
where $e_{1}=[1,0,\cdots,0]^{\top}\in\mathbb{R}^{N_{\textrm{bus}}}$, $R=\textrm{{diag}}(\textrm{r})$, $X=\textrm{{diag}}(\textrm{x})$. Note that the $\alpha$-th element in vector $\mathbf{r}$ (resp. $\mathbf{x}$) denotes the resistance (resp.) reactance of branch $\alpha$, $U_{k}$ denotes the vector of squared voltage magnitude at the time instant $k$, $p^{\textrm{pcc}}_{k}$, $q^{\textrm{pcc}}_{k}$ denote active and reactive power exchanges with the transmission system at the PCC of the distribution system. We use vectors $P^{d}_{k}$, $Q^{d}_{k}$ to denote the active and reactive nodal power consumptions, $P^{l}_{k}$, $Q^{l}_{k}$ to denote the active and reactive branch flows, and $P^{s}_{k}$ to denote the nodal consumptions by distributed energy storage systems (ESS)at time period $k$. Moreover, (1a) fixes the voltage magnitude at the slack bus. Equations (1b)-(1d) are the LinDistFlow constraints. Upper and lower bounds (1f) restrict the voltage magnitude at each bus and the charging/discharging power of ESS s. We rewrite the above power flow equations (1a)-(1d) in a compact form:
$M\chi_{k}+BP^{s}_{k}+D_{k}=0,$ (2)
where
$M$ $\,=\,\begin{bmatrix}e_{1}^{\top}&0&0&0&0\\ C^{\textrm{inc}}&-2R&-2X&0&0\\ 0&-(C^{\textrm{inc}})^{\top}&0&e_{1}&0&\\ 0&0&-(C^{\textrm{inc}})^{\top}&0&e_{1}\end{bmatrix},$
$\chi_{k}$ $\,=\,\begin{bmatrix}U\\ P^{l}\\ Q^{l}\\ p^{\textrm{pcc}}\\ q^{\textrm{pcc}}\end{bmatrix},\;B=-\begin{bmatrix}0\\ 0\\ C^{s}\\ 0\end{bmatrix},\;\text{and}\;\;D_{k}=\begin{bmatrix}e_{1}\\ 0\\ P^{d}_{k}\\ Q^{d}_{k}\end{bmatrix}.$
Note that $M\in\mathbb{R}^{(2N_{\textrm{bus}}+N_{\textrm{line}}+1)\times(N_{\textrm{bus}}+2N_{\textrm{line}}+2)}$ is a square matrix since for radial distribution grids, $N_{\textrm{bus}}=N_{\textrm{line}}+1$. In (2), $M$ and $B$ remain time-invariant. All dependent variables $\chi_{k}$ are influenced by controllable power injections $P^{s}_{k}$ from ESS s, as well as the load demands $P_{k}^{d}$ and $Q_{k}^{d}$ at each time period $k$. Therefore, in this paper, the flexibility in distribution systems primarily arises from the integration of ESS s.
In [20], it is shown that the squared voltage magnitude $U$ can be explicitly expressed by the active and reactive power injections. However, positive definiteness of resistance and reactance for all the branches is required, a condition not universally met in practical power system datasets, as discussed in [21]. To extend the applicability of the proposed coordination framework to a broader range of power systems, we generalize the result from [20]. With the assistance of graph theory, we rigorously demonstrate the invertibility of matrix $M$, affirming that all state variables, including squared voltage magnitude $U$, can be explicitly expressed in terms of controllable power injections for all radial networks. This expansion significantly enhances the robustness and versatility of the proposed framework for practical power systems.
Lemma 1.
[…]
This lemma follows directly from the fact that a leaf has only one parent in a radial network.
Lemma 2 ([19]).
A radial network with at least two buses has at least two leaves. Deleting a leaf from a radial network with $N$ buses produces a radial network with $N-1$ buses.
Proposition 1.
Given a radial network $\mathcal{G}$, matrix $M$ is invertible.
The proof can be found in the Appendix. As a result of the generalized proposition 1, for a given distribution grid, all the dependent variables in $\chi_{k}$ can be expressed explicitly in terms of the controllable power injections $P^{s}_{k}$, and thus, the exact feasible set is convex and can be written in a convex polytope with respect to $P^{s}_{k}$,
$\mathcal{P}_{k}^{s}=\{P^{s}_{k}\in\mathbb{R}^{e}|A^{s}P^{s}_{k}\leq b^{s}\},$ (3)
where $e$ denotes the number of ESS s. In the example illustrated in Fig. 2, the blue polytope represents an exact feasible set constrained by upper and lower voltage bounds along with power limits of ESS s (1f).
II-A2 Maximum Volume Inner Approximation
In this paper, the flexibility of distribution systems primarily arises from the integration of ESS in (II-A1). Instead of applying the exact feasible set (3), we replace the complex polytope with a strictly inner hyperbox approximation, enhancing computational efficiency while maintaining safe operation guarantees within the system, i.e.,
$\mathcal{B}^{s}_{k}\subseteq\mathcal{P}^{s}_{k},\quad\forall k\in\{1,2,\cdots,N_{k}\},$ (4)
where the hyperbox $\mathcal{B}_{k}$ is defined as
$\mathcal{B}^{s}_{k}(\underline{P}^{\textrm{appr}}_{k},\overline{P}^{\textrm{appr}}_{k})=\{P^{s}_{k}\in\mathbb{R}^{e}|\underline{P}^{\textrm{appr}}_{k}\leq P^{s}_{k}\leq\overline{P}^{\textrm{appr}}_{k}\},$ (5)
Note that $\overline{P}^{\textrm{appr}}_{k}$ and $\underline{P}^{\textrm{appr}}_{k}$ are upper and lower bounds of the inner hyperbox approximation. To maximize the performance of the resulting ESS system, we adopt the so-called maximum volume inner hyperbox [22]. The hyperbox (5) can be written as $\mathcal{B}^{s}_{k}(\zeta,\zeta+\xi)$ and the inner approximation can be obtained by solving the following optimization problem:
$\max_{\xi,\zeta}$ $\quad\sum_{i\in\mathcal{E}}\ln\zeta_{i},$ (6a)
s.t. $\quad A^{s}\xi+A^{s+}\zeta\leq b^{s},$ (6b)
where $A^{s+}$ is the positive part of $A^{s}$ and =$\mathcal{E}=\{1,\cdots,e\}$ is the set of ESS s. However, in practice, it can occur that the standby mode of a ESS is excluded by the inner approximation, i.e.,
$\exists i\in\mathcal{E},\quad[\overline{P}^{\textrm{appr}}_{k}]_{i}<0\quad\textrm{or}\quad[\underline{P}^{\textrm{appr}}_{k}]_{i}>0,$
i.e., the origin is not included in the resulting hyperbox (green), as shown in Fig. 2 (a).
To address this issue, instead of focusing on maximizing the volume in $\mathbb{R}^{e}$ space, i.e., finding an equilibrium where ESS s have wide ranges of permissible power output intervals, we propose to maximize the volume within the $\mathbb{R}^{2d}$ space, thereby expanding both charging and discharging power limits of the ESS s according to
$\max_{\xi,\zeta}$ $\quad\sum_{i\in\mathcal{E}}\ln(\xi_{i}+\zeta_{i})+\ln(-\xi_{i}),$ (7a)
s.t. $\quad A^{s}\xi+A^{s+}\zeta\leq b^{s},$ (7b)
enabling scenarios where, for instance, both ESS s can charge even during periods of high system load, as illustrated in Fig. 2(b), but more importantly, the origin is included in the hyperbox (green).
(a) $\mathcal{B}_{k}^{s}$ calculated by (II-A2)
(b) $\mathcal{B}_{k}^{s}$ calculated by (II-A2)
Figure 2: Comparison of inner approximation methods with 2 ESS s located in the heavily loaded IEEE 33-bus system. The orange and the purple lines show the upper and lower bounds on the squares of voltage magnitudes (1e). The red lines show the limits on ESS s’ power output (1f). The blue polytope denotes the exact feasible set (3) and the green rectangle denotes the inner hyperbox approximation (5).
II-B Coordinated Economic Dispatch for ITD Systems
II-B1 Aggregating Distribution Systems
Since LinDistFlow (II-A1) ignores the power losses along branches, power exchanged with the transmission can be expressed with the assistance of conservation of power for all time periods $k\in\mathcal{K}$,
$p_{k}^{\textrm{pcc}}=$ $\mathds{1}^{\top}P_{k}^{d}+\mathds{1}^{\top}P_{k}^{s},\quad P_{k}^{s}\in\mathcal{B}_{k}^{s},$ (8a)
$q_{k}^{\textrm{pcc}}=$ $\mathds{1}^{\top}Q_{k}^{d},$ (8b)
where $\mathcal{B}_{k}^{s}$ is calculated by applying (II-A2). As a result, at the transmission level, a distribution system can be modeled as a load and multiple ESS s at the PCC.
II-B2 Multiperiod AC Optimal Power Flow
We utilize MPOPF for coordinating different TSOs at the transmission level, i.e.,
$\min\quad$ $\sum_{k=1}^{N_{k}}(P^{g}_{k})^{\top}\textrm{{diag}}(a_{2})P^{g}_{k}+a_{1}^{\top}P^{g}_{k}+a_{0}^{\top}\mathds{1},$ (9a)
subject to $\forall k\in\mathcal{K}:=\{1,2,\cdots,N_{k}\}$
$P^{b}_{k}(\Theta_{k},V_{k})$ $=C^{g}P^{g}_{k}-P^{d}_{k}-{\color[rgb]{0,0,0}C^{\textrm{pcc}}P^{\textrm{pcc}}_{k}}-C^{s}P^{s}_{k},$ (9b)
$Q^{b}_{k}(\Theta_{k},V_{k})$ $=C^{g}Q^{g}_{k}-Q^{d}_{k}-{\color[rgb]{0,0,0}C^{\textrm{pcc}}Q^{\textrm{pcc}}_{k}},$ (9c)
$\left|S_{k}^{l}(\Theta_{k},V_{k})\right|$ $\leq\overline{S}^{l},$ (9d)
$\underline{V}\leq V_{k}\leq\overline{V}$ $,\,\underline{P}^{g}\leq P^{g}_{k}\leq\overline{P}^{g},\,\underline{Q}^{g}\leq Q_{k}^{g}\leq\overline{Q}^{g},$ (9e)
$E_{k}=E_{k-1}+$ $\Delta t\cdot P^{s}_{k}\;\text{ with initial state }E_{0}=E(t),$ (9f)
$P^{g}_{k}=P^{g}_{k-1}+$ $\Delta P^{g}_{k}\;\text{ with initial state }P_{0}^{g}=P_{0}^{g}(t),$ (9g)
$\underline{E}\leq E_{k}\leq\overline{E}$ $,\,\underline{P}_{k}^{s}\leq P_{k}^{s}\leq\overline{P}_{k}^{s},\,\underline{R}\leq\Delta P_{k}^{g}\leq\overline{R},$ (9h)
where $P^{b}_{k}$, $Q^{b}_{k}$ represent the vectors of active and reactive power injections for all buses at time period $k$, expressed as nonlinear mappings $\mathbb{R}^{N_{\textrm{bus}}}\times\mathbb{R}^{N_{\textrm{bus}}}\mapsto\mathbb{R}^{N_{\textrm{bus}}}$ with respect to nodal voltage angles $\Theta_{k}$ and magnitudes $V_{k}$. Similarly, $S^{l}_{k}$ are nonlinear mappings $\mathbb{R}^{N_{\textrm{bus}}}\times\mathbb{R}^{N_{\textrm{bus}}}\mapsto\mathbb{R}^{N_{\textrm{line}}}$ representing apparent branch power flows for all branches at time period $k$; for the detailed formulation of power injections and branch power flows, we refer readers to [23]. $P^{\textrm{pcc}}_{k}$, $Q^{\textrm{pcc}}_{k}$ stack active and reactive power exchanges (II-B1) for all distribution systems, respectively. Evidently, MPOPF (II-B2) constructs a simultaneous formulation of $N_{k}$ AC OPF problems with standard power flow constraints (9b)-(9e), coupled by intertemporal interactions (9f) (9g) and the corresponding upper and lower bounds (9h). Notice that ESS s possess time-variant power limits in (9h), due to the inner hyperbox approximation (5) utilized for aggregating distribution systems.
Rather than devising intricate mathematical models to precisely represent distribution systems, the flexibility aggregation method offers a substantial reduction in computational complexity of the MPOPF in (II-B2). It enhances the scalability of the proposed framework without sacrificing the active involvement of distribution systems in the dispatch problems.
III Methodology
This section presents the proposed distributed real-time coordination framework of ITD systems using a receding horizon scheme while considering day-ahead forecast and actual values. Compared to the classical distributed MPC scheme, only solving the structured optimal control problem either in a spatially distributed manner or in a temporally distributed manner, our approach decouples the optimization problems across both different system operators and periods, with each subproblem representing an individual single-period AC OPF of a single transmission system.
III-A Distributed Formulation
We describe a coordination problem of ITD systems by a tuple $\mathcal{C}=(\mathcal{N},\,\mathcal{L},\,\mathcal{K},\,\mathcal{R})$ over $N_{k}$ time periods. Thereby, $\mathcal{N}$ denotes the set of all buses, $\mathcal{L}$ the set of all branches, and =$\mathcal{R}=\{T_{1},\,T_{2},\,\cdots\}$ denotes the set of coordinated transmission systems.
The objective function (9a) summarizes quadratic generation cost from all regions $\ell\in\mathcal{R}$ over all time periods $k\in\mathcal{K}$. This enables a straightforward separation of the objective function across different system operators and time periods:
$f(x)=\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{R}}f_{k,\ell}(x_{k,\ell}),$
where $x_{k,\ell}$ represents state variables in the transmission system $\ell$ at the time period $k$ and $x$ is a vector stacking all the subvectors $x_{k,\ell}$.
The constraints of MPOPF (II-B2) can be decoupled across time periods, where each of the temporal coupling constraints (9f) (9g) is associated with only one specific transmission system. Thereby, these temporal coupling constraints can be written in the following standard affinely coupled form
$\sum_{k\in\mathcal{K}}\Lambda_{k,\ell}x_{k,\ell}={\color[rgb]{0,0,0}0},\;\ell\in\mathcal{R},$
where the sparse matrices $\Lambda_{k,\ell}$ contain non-zero elements of $\{-1,1,\Delta t\}$, connecting the state variables current $E_{k,\ell}$ and $P^{g}_{k,\ell}$ with neighboring time periods $\{k-1,k+1\}$ for each transmission system $\ell$. Regarding spatial coupling among different TSOs, we follow the idea of sharing components [24], i.e., sharing nodal voltage angles and magnitudes at both sides of connecting tie-lines between neighboring transmission systems. The resulting spatial coupling constraints are linear and can be written in the following affinely coupled form
$\sum_{\ell\in\mathcal{R}}\Gamma_{k,\ell}x_{k,\ell}={\color[rgb]{0,0,0}0},\;k\in\mathcal{K},$
where the sparse matrices $\Gamma_{k,\ell}$ contain non-zeros elements of $\{-1,1\}$, connecting the coupling voltage angles and magnitudes between neighboring transmission systems for each time period $k$.
Thereby, the MPOPF in the transmission level can be decoupled across different system operators and time periods and reformulated in standard affinely distributed form
$\min\;$ $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{R}}f_{k,\ell}(x_{k,\ell})$ $\begin{array}[]{c}\textsc{Decoupled}\\ \textsc{Objective}\end{array}$ (10c)
s.t. $\forall k\in\mathcal{K},\;\sum_{\ell\in\mathcal{R}}\Gamma_{k,\ell}x_{k,\ell}=0\;\,\mid\lambda_{k}$ $\begin{array}[]{c}\textsc{Spatial}\\ \textsc{Couplings}\end{array}$ (10f)
s.t. $\forall\ell\in\mathcal{R},\;\sum_{k\in\mathcal{K}}\Lambda_{k,\ell}x_{k,\ell}=0\;\,\mid\kappa_{\ell}$ $\begin{array}[]{c}\textsc{Temporal}\\ \textsc{Couplings}\end{array}$ (10i)
[…]
where $\lambda_{k}$, $\kappa_{\ell}$ and $\nu_{k,\ell}$ denote Lagrangian multipliers associated with the corresponding constraints. Constraints (10l) correspond to the standard AC OPF constraints (9b)-(9e) with power and energy limits on the ESS s (9h) for each transmission system $\ell\in\mathcal{R}$ over all time periods $k\in\mathcal{K}$.
III-B Real-Time Distributed Coordination Scheme
As (II-B2) is reformulated in an affine-coupled distributed form (III-A), it can be solved efficiently by using distributed optimization algorithms. In this paper, we tailor the ALADIN algorithm [14] to deal with (III-A) in a closed loop. The resulting distributed coordination scheme in receding horizon fashion is outlined in Algorithm 1.
Based on the local measurements collected in Step 1), Step 2) of Algorithm 1 outlines the tailored ALADIN algorithm to solve (III-A). Step 2.a) solves $N_{k}\cdot|\mathcal{R}|$ subproblems, in which the regional TSO deals with the $N_{k}$ temporal subproblems in parallel locally. These problems are constructed using the Lagrangian of (III-A) by dualizing the spatial coupling (10f) and temporal coupling (10i). Based on the decoupled solutions $y_{k,\ell}$, Step 2.b) computes sensitivities of objective and constraints with respect to the current iteration of ALADIN. Here, in order to improve the numerical robustness of the algorithm, a small perturbation is added to the second-order derivatives (13) approximated by positive definite $H_{k,\ell}$ Notice that under a mild assumption for the perturbation as outlined in [3, Theorem 2], the local quadratic convergence can be guaranteed. Step 2.c) solves the coupled QP (2c) with only equality constraints. Taking the temporal coupling (14c) as local equality constraints for the $\ell$-th region, one can solve (2c) in a decentralized manner that only requires neighbor-to-neighbor communications. For more details, the reader is referred to [25]. A practical stopping criterion is to stop if the following conditions hold,
$\max_{k\in\mathcal{K}}\left\|\sum_{\ell\in\mathcal{R}}\Gamma_{k,\ell}y_{k,\ell}\right\|\leq\epsilon,\;\max_{\ell\in\mathcal{R}}\left\|\sum_{k\in\mathcal{K}}\Lambda_{k,\ell}y_{k,\ell}\right\|\leq\epsilon.$
After local solutions are determined at the transmission level, the inputs $(P_{0,\ell}^{s,*},\Delta P_{0,\ell}^{g,*})$ are deployed to the respective generators and storages. Step 4) in Algorithm 1 is a warm-start for Step 2) for the next online iteration following the strategy proposed in [26].
Algorithm 1 Distributed Real-Time Coordination of ITD Systems
Offline:
Repeat:
IV Case Study
For an illustrative case study, copies of the 118 bus system from the PGLib-OPF [27] are used for the transmission systems and the IEEE 33 bus system for the distribution systems. Four transmission systems are interconnected through multiple connecting tie-lines, as shown in Fig. 1(b), with each connected to 10 distribution systems, roughly representing the 4 TSOs as in Germany. Consequently, the ITD system comprises a total of $1792$ buses. By applying the proposed flexibility model of distribution networks, the resulting MPOPF problems, in the end, will solve dispatch schedules for a system having $472$ buses at the transmission level. Simulations are run for 24 hours with a prediction horizon of $N_{k}=96$ and a time interval of $\Delta t=15$ min. The resulting optimization problems in the transmission level have $187,776$ state variables partitioned into $384$ subproblems. To simulate modern contemporary power system scenarios, we utilize measurement data from the ENTSO-E Transparency Platform [17] dated July 24, 2023 [28]. This day was characterized by heavy rainfall and stood out for relatively significant large prediction mismatches, particularly in solar generation. The dataset comprises day-ahead predictions and actual values for load demand, solar generation, and wind generation, as shown in Fig. 3.
Figure 3: Day-head forecasts and actual values of load demand, solar generation, and wind generation for 4 TSOs in Germany from ENTSO-E platform [17]
Figure 4: Optimal economic dispatch by isolated approach for 4 TSOs during simulation
Figure 5: Optimal economic dispatch by coordinated approach for 4 TSOs during simulation
Figure 6: State-of-Charge by coordinated approach for 4 TSOs during simulation
Figure 7: Convergence Performance of distributed NMPC
In the case study, three distinct operating strategies are introduced and analyzed, i.e., isolated operation, centralized coordination, and distributed coordination. In all these strategies, the flexibility of distribution systems is aggregated into transmission systems as proposed in Sec. II. In the isolated operation mode, transmission systems are operated independently without communication or power sharing among different transmission systems. The results of the economic dispatches per time period are comprehensively visualized in Fig. 4.
Centralized and distributed coordinations employ the system model in (II-B2), treating all systems as a single grid at the transmission level. Unlike isolated operation, the coordinated operating modes facilitate power exchange between transmission systems to minimize overall generation costs, as illustrated in Fig. 5. The difference lies in the optimization approaches. Centralized coordination communicates all private data to a centralized entity and employs a centralized approach to solve the optimization problem. In contrast, distributed coordination solves the optimization problem based on the proposed algorithm in a distributed fashion with limited information exchanged.
A noteworthy instance occurs at $13:00$ where transmission system $T_{1}$ encounters a significant prediction mismatch. In this time period, $T_{1}$ experiences higher actual demand and reduced solar generation, coinciding with lower SOC of ESS s in $T_{1}$, as visualized in Fig. 3 and Fig. 6. In response to this prediction error, power export to other systems is intentionally curtailed as a compensatory measure.
A comparison of economic efficiency is presented in Table I. Notably but not surprisingly, isolated operation stands out with the highest total cost. In contrast, centralized coordination can attain the lowest overall cost. Meanwhile, distributed coordination offers an intriguing alternative, preserving data privacy with limited information exchange while achieving competitive overall costs. Both coordination approaches can find a local minimizer of the system model (II-B2), and optimality gaps in terms of total costs are negligible.
The detailed convergence performance of the proposed distributed approach over $24$ hours is demonstrated in Fig. 7, with the centralized solutions employed as reference solutions. Notably, all the 96 optimization problems in the daily operation exhibit fast convergence in a dozen iterations, within 500 seconds, while both the primal and dual residuals converge to tolerable values. Compared with centralized coordination, the proposed distributed approach showcases remarkable accuracy in terms of controllable power injections and total cost gap over all 96 time periods. These results highlight the scalability and numerical robustness for real-world applications in large-scale ITD systems.
TABLE I: Generation Costs [] with aggregated Flexibility of DSOs
Isolated Centralized Distributed
Coordination Coordination
$T_{1}$ 2 034 05220340522\,034\,0522 034 052 2 499 73624997362\,499\,7362 499 736 2 499 78524997852\,499\,7852 499 785
$T_{2}$ 2 006 14520061452\,006\,1452 006 145 2 396 54923965492\,396\,5492 396 549 2 396 57323965732\,396\,5732 396 573
$T_{3}$ 5 597 56655975665\,597\,5665 597 566 4 058 84640588464\,058\,8464 058 846 4 058 84240588424\,058\,8424 058 842
$T_{4}$ 1 778 17917781791\,778\,1791 778 179 2 241 37622413762\,241\,3762 241 376 2 241 36822413682\,241\,3682 241 368
Total 11 415 9421141594211\,415\,94211 415 942 11 196 5051119650511\,196\,50511 196 505 11 196 5681119656811\,196\,56811 196 568
V Conclusion
This paper proposes a distributed framework designed for efficient coordination of ITD systems. Within the framework, the flexibility inherent in distribution systems is aggregated and provided to the respective transmission systems, and the coordination of TSOs is determined by employing an aladin-type distributed nmpc scheme, where the resulting optimization problems are divided across different TSOs and time periods. Numerical simulations conducted on real operational data with relatively large prediction mismatches indicate our approach’s practical applicability in real-world scenarios.
Considering a non-slack leaf bus $\beta$, $(\alpha,\beta)$ is the only nonzero element in $\beta$-th column in matrix $C^{\textrm{inc}}$ due to the incidence matrix property in Lemma 1. Hence, $(\alpha+1,\beta)$ is the only nonzero element in $\beta$-th column in matrix $M$. Similarly, $(\beta+N^{\textrm{bus}},\beta+N^{\textrm{bus}})$ and $(\beta+2N^{\textrm{bus}},\beta+2N^{\textrm{bus}})$ are the only nonzero elements in the $(\beta+N^{\textrm{bus}})$-th and the $(\beta+2N^{\textrm{bus}})$-th row respectively.
By eliminating the leaf bus $\beta$ of the network $\mathcal{G}$, we obtain a reduced radial network $\mathcal{G}^{(1)}$. The resulting matrix $M^{(1)}$ can be viewed as a submatrix of $M$ by removing the set of row $\{\alpha+1,\beta+N^{\textrm{bus}},\beta+2N^{\textrm{bus}}\}$ and the set of column $\{\beta,\alpha+N^{\textrm{bus}},\alpha+2N^{\textrm{bus}}\}$.
Since the nonzero elements in the incidence matrix $C^{\textrm{inc}}(\mathcal{G})$ is {1, $,\,1\}$ 1 }, the determinant of matrix $M$ can be written as
$\left|\textrm{det}(M)\right|=\left|\textrm{det}(M^{(1)})\right|$ (15)
with the assistance of cofactor expansions.
By further removing non-slack leaves of the resulting reduced radial networks, we have
$\left|\textrm{det}(M)\right|=$ $\left|\textrm{det}(M^{(1)})\right|=\cdots=\left|\textrm{det}(M^{(N^{\textrm{bus}}-1)})\right|$
$=$ $\textrm{det}\left(\begin{bmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{bmatrix}\right)=1.$ (19)
Therefore, $M$ is invertible for the given radial network.